OpenStudy (anonymous):

log3^(3x-5)=log3^(4x+4)

OpenStudy (anonymous):

3x-5=4x+4 x=9

OpenStudy (anonymous):

I MEAN -9

OpenStudy (anonymous):

log x^n = n logx so log3^(3x-5) = (3x-5)log3 Threfore, (3x-5)log3=(4x+4) log3 so 3x-5=4x+4 x=-9

OpenStudy (anonymous):

This is from a test I took. I got -9 also, but it was marked wrong.

OpenStudy (anonymous):

Time to kick some teacher's retricethen

OpenStudy (anonymous):

lol retrice, i mean a$$

OpenStudy (anonymous):

LOL